### v2

# Routoo: Learning to Route to Large Language Models Effectively

Abstract
Developing foundational large language models (LLMs) is becoming increasingly costly and inefficient. Also, closed-source and larger open-source models generally offer better response quality but come with higher inference costs than smaller models. In this paper, we introduce Routoo, an architecture designed to optimize the selection of LLMs for specific prompts based on performance, cost, and efficiency. Routoo consists of two key components: a performance predictor and a cost-aware decoding. The performance predictor is a lightweight LLM that estimates the performance of various underlying LLMs without needing to execute and evaluate them. The cost-aware decoding then selects the most suitable model based on these predictions and other constraints like cost and latency.
We evaluated Routoo using the MMLU benchmark across 57 domains employing open-source models. Our results show that Routoo matches the performance of the Mixtral 8x7b model while reducing inference costs by one-third. Additionally, by allowing increased costs, Routoo surpasses Mixtral’s accuracy by over 5% at equivalent costs, achieving an accuracy of 75.9%. When integrating GPT4 into our model pool, Routoo nearly matches GPT4’s performance at half the cost and exceeds it with a 25% cost reduction. These outcomes highlight Routoo’s potential to create new SOTA in a cost-effective manner by leveraging the collective knowledge of multiple LLMs.
1 Introduction
Developing foundational models OpenAI et al. (2024); Touvron et al. (2023); Jiang et al. (2023a, 2024) is capital-intensive, necessitating vast computational resources and extensive, high-quality data Minaee et al. (2024). Furthermore, the field is nearing the upper bounds of network size and data capacity, resulting in progressively marginal enhancements over existing models Udandarao et al. (2024). This scenario echoes a critical juncture in human advancement, where the ’divide and conquer’ methodology emerges as a viable and scalable alternative. This approach entails cultivating domain-specific experts and judiciously harnessing them to forge a composite, high-performance model.
The capabilities of existing LLMs appear to be complementary to a significant degree. An illustrative case is the MMLU benchmark Hendrycks et al. (2021), where selecting the optimal open-source model for each question hypothetically yields an accuracy of 97.5%, at a computational cost akin to a model with 13 billion parameters.  [^1] Further details are provided in Appendix A. In contrast, GPT4 OpenAI et al. (2024) achieves an accuracy of 86.4%, while Mixtral 8x7b Jiang et al. (2024), as the leading open-source model, reaches 70%. These figures suggest considerable scope for integrating LLMs’ knowledge to create new state-of-the-art models. Additionally, it is noteworthy that many practical tasks do not require intricate reasoning and can be efficiently addressed by models of moderate complexity; only a minority of tasks demand the advanced capabilities of a model like GPT4 Zaharia et al. (2024); Ding et al. (2024); Shnitzer et al. (2023); Chen et al. (2023). The landscape of open-source models, with approximately 450,000 entries on Hugging Face, [^2] https://huggingface.co is both dynamic and expansive. The rapid proliferation of new models, particularly smaller, domain-focused ones, poses a challenge in tracking and leveraging the latest advancements effectively.
Figure 1: Routoo architecture has two main components: a performance predictor and a cost-aware decoder. The former predicts the accuracy of experts for a specified query, and the latter chooses the underlying LLM and generates the response by considering the cost and efficiency of each model. The Universe Constructor builds a complementary set of underlying models.
We propose Routoo, an architecture to learn to leverage a universe of trained LLMs automatically and dynamically to create a better-performing model.
The cornerstone of Routoo is a lightweight LLM that predicts the performance of each underlying LLM on queries without the need for actual execution and evaluation. Based on these predictions, a cost-aware decoding algorithm selects the most suitable model, prioritizing factors such as inference cost and speed. For instance, when faced with a task that is predicted to be performed nearly equally well by a 7 billion parameter model or a more extensive 70 billion parameter model, the Routoo will opt for the former when factors like speed and cost-efficiency are prioritized. This approach ensures optimal resource utilization without compromising on quality.
Our architecture marks a significant departure from traditional Mixture of Experts (MoE) models Lieber et al. (2024); Jiang et al. (2024); Fedus et al. (2021); Shazeer et al. (2017). While MoE relies on gating over various expert sub-networks within each layer to predict the next token, it requires all expert parameters to be loaded onto a single, high-end machine. This limitation hinders scalability in the number of experts. In contrast, each ’expert’ within our system operates independently and can be hosted on different machines, potentially utilizing a different neural network architecture. This flexibility means we can incorporate a vast array of experts, ranging from those specializing in system-level Java programming to those adept at curating travel experiences in London. Moreover, Routoo can be tuned to leverage only smaller models, which is particularly beneficial for users without access to high-end computing resources (e.g., GPUs with 100GB). Furthermore, our architecture facilitates easy integration of additional optimization criteria, including cost, speed, and privacy considerations.
We evaluate our Routoo on MMLU benchmark Hendrycks et al. (2021). We show that it achieves competitive performance with Mixtral 8x7b Jiang et al. (2024) while only consuming two-thirds of its inference cost. Increasing the cost budget allows Routoo to outperform Mixtral by 5% at the same cost level, reaching an accuracy of 75.9%. By adding GPT4 OpenAI et al. (2024) as one of the underlying experts, our Routoo achieves competitive performance with GPT4 at half the cost and exceeds it with a 25% cost reduction.
To summarize, our contributions are:
  • We propose Routoo, an LLM-based system designed to intelligently identify the best-performing LLM for a given query while considering constraints (e.g. cost), effectively integrating knowledge of multiple LLMs to create a new state-of-the-art model.
  • We evaluate our architecture on MMLU benchmark, and significantly outperform Mixtral 7x8b by 5% with similar inference cost. Also, our Routoo achieves competitive performance with GPT4 at half the inference cost and surpasses it by reducing the inference cost by 25%.
2 Related Work
LLM Benchmarks.
In recent years, many benchmarks Wang et al. (2024); Liang et al. (2023); Beeching et al. (2023); Gao et al. (2021); Hendrycks et al. (2021) are created to identify the capability of LLMs. The Massive Multitask Language Understanding  (MMLU) benchmark Hendrycks et al. (2021) contains multiple-choice question-answer (MCQA) pairs from various branches of knowledge. This includes 57 domains, including mathematics, US history, computer science, and law. OpenLLM leaderboard Beeching et al. (2023) includes several question-answering datasets, e.g. HellaSwag Zellers et al. (2019), Truthfulqa Lin et al. (2022), Winogrande Sakaguchi et al. (2019), GSM8k Cobbe et al. (2021), ARC Clark et al. (2018) and MMLU, which is used to compare open-source LLMs. HELM Liang et al. (2023) consists of 42 scenarios covering a variety of use cases. In this work, we utilized the MMLU benchmark Hendrycks et al. (2021) to assess our models. This benchmark spans several domains and is recognized as one of the challenging benchmarks on the OpenLLM leaderboard Lu et al. (2023).
Model Selection.
Previous works in selecting the best LLMs mainly focus on identifying the one that generates the most optimal output for a given input. Liu and Liu (2021); Ravaut et al. (2022) proposed specialized scoring or re-ranking models that can be used for the generation tasks (summarisation, here). Speculative decoding Kim et al. (2023); Leviathan et al. (2023) accelerates the decoding of expensive models by using small, efficient decoders for the ’easy’ steps. This approach can complement routing methods. LLM-BLENDER Jiang et al. (2023b) is an ensembling framework to reach better performance by mixing the results of LLMs with a ranking module, followed by a generation module to generate from top candidates of the ranker. FrugalGPT Chen et al. (2023) executes experts sequentially until an expert reaches the acceptable generation performance. HybridLLM Ding et al. (2024) proposes a routing approach to direct queries to the suitable expert, utilizing two underlying LLMs. Different from previous work, our Routoo identifies the most suitable expert without executing the inference of underlying LLMs. Additionally, our approach can be efficiently generalized to scenarios involving many underlying LLMs.
3 Architecture
3.1 Problem Formulation
Given a set of $N$ LLM $e_{1},e_{2},\ldots,e_{N}$ and a set of $m$ queries $q_{1},q_{2},\ldots,q_{m}$, Our goal is to assign the most cost-effective model $e_{i}$ to each query $q_{j}$that can accurately answer the query.
For each model-query pair, we define an evaluation function $\operatorname{eval}(e_{i}(q_{j}))$ [^3] Evaluation function can have the reference as an optional input., which returns a score from 0 to $K-1$ indicating the correctness of the model’s response, where 0 is unacceptable and $K-1$ is optimal. Each response also incurs a cost $\operatorname{cost}(e_{i}(q_{j}))$.
The objective is to maximize the total correctness scores across all queries while keeping the total cost within a budget $B$:
$\max_{\pi}$ $\ \sum_{j=1}^{m}\operatorname{eval}(\pi(q_{j})(q_{j}))$
s.t. $\ \sum_{j=1}^{m}\operatorname{cost}(\pi(q_{j})(q_{j}))\leq B$
Here, $\pi$ is the function that assigns each query to a model, based on both cost and correctness. In this formulation, the cost can be multi-faceted, encompassing aspects like computational cost, speed, etc. However, for simplicity, we consider a singular budget constraint $B$ in this context. To solve this without exhaustively testing every model on every query—a prohibitive approach—we divide the problem into two manageable parts:
  1. 1.
    Performance Prediction: This component approximates the evaluation score $\operatorname{eval}(e_{i}(q_{j}))$ for each model on each query without actual execution.
  2. 2.
    Cost-Aware Decoder: This step selects the best model for each query, balancing accuracy and cost-effectiveness using the performance predictions from the previous component.
These components work together to dynamically and efficiently allocate models to queries, as illustrated in Figure 1. In the following sections, we will delve deeper into each component.
3.2 Performance Predictor
Figure 2: The performance predictor: Given the query and an LLM, it outputs the performance prediction $\operatorname{\hat{eval}}(e_{i},q_{j})$.
The performance predictor is a lightweight LLM designed to estimate the effectiveness of each underlying LLM for a given query. It predicts the evaluation score, representing the expected response quality of a model to a query. This prediction is formulated as:
$\operatorname{\hat{eval}}(e_{i},q_{j})=\operatorname{Pred}(e_{i},q_{j}),$ (1)
where $\operatorname{\hat{eval}}(e_{i},q_{j})$ denotes the predicted score of model $e_{i}$ for query $q_{j}$, and $\operatorname{Pred}(e_{i},q_{j})$ is the predictive model.
The process, illustrated in Figure 2, involves:
$\begin{split}\begin{cases}h_{q_{j}}=\operatorname{Enc}(q_{j})\\ h_{e_{i}}=\operatorname{Emb}(e_{i})\\ s_{q_{j},e_{i}}=\operatorname{Linear}(h_{q_{j}}-h_{e_{i}})\end{cases}\end{split}$
where $\operatorname{Enc}(.)$ is the encoder of the input query. Inspired by Radford et al. (2019), we use a decoder-only model as the encoder and extract the embedding of the last token as the representation of the input query. $Emb(.)$ is an embedding layer, where each embedding is assigned to a specific model. $Linear(.)$ is a linear layer to convert the combination of query and model embeddings to the logits of the evaluation score. $s_{q_{i},e_{j}}$ is a list of predicted evaluation scores for a given model and query. The size of this list depends on the number of levels (K) in the evaluation score e.g. 2 for binary evaluation scores.
The model is trained by minimising the cross-entropy loss function Good (1952) of $s_{q_{i},e_{j}}$ and $\operatorname{eval}(e_{j}(q_{i}))$ over all $N$ expert models, and $m$ queries as:
$\operatorname{min}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{N}\operatorname{CE}(s_{q_{i},e_{j}},\operatorname{eval}(e_{j}(q_{i})))$ (2)
where $\operatorname{CE}(.)$ is the cross-entropy loss.
During the inference time, $\operatorname{argmax}(.)$ is applied for a given model and query as:
$\operatorname{\hat{eval}}(e_{i},q_{j})=\operatorname{argmax}_{(0,...,K-1)}(s_{q_{j},e_{i}})$ (3)
where $\operatorname{argmax}(.)$ is applied over the levels of the evaluation score.
3.3 Cost-Aware Decoder
The second phase of our architecture involves the decoding step, where the estimated scores from the performance predictor are used to determine the optimal assignment of models to queries. Our objective is to maximize the overall effectiveness of the responses within a given budget constraint. The optimization problem is formulated as follows:
$\max_{\pi}$ $\ \sum_{j=1}^{m}\operatorname{\hat{eval}}(\pi(q_{j}),q_{j})$
s.t. $\ \sum_{j=1}^{m}\operatorname{cost}(\pi(q_{j})(q_{j}))\leq B$
where $\pi(.)$ determines the expert assignments and $B$ represents the predefined budget constraint.
We propose a greedy algorithm to approximate this optimization. Noting the challenges of accurately estimating the exact cost of a model for a specific query, instead, we consider the average cost $c_{i}$ for a model $e_{i}$ responding to an average length input and response.
Result: Find the set U of size k that maximizes $S(U)$ according to equation (4)
Initialize an empty set $U=\{\}$
Set budget to 0
while budget < k do
       1. Identify the model $e_{i}^{*}$ that, when added to $U$ forming $U^{*}$, maximizes $S(U^{*})$
       2. Update $U$ to $U^{*}$
       3. Increment budget by 1
       if No further improvement in $S(U)$ then
             break;
       end if
end while
Algorithm 1 Universe Construction.
The algorithm includes the following steps:
  1. 1.
    Performance-to-Cost Ratio: For each query $q_{j}$ and model $e_{i}$, calculate the performance-to-cost ratio. Let $\operatorname{\hat{eval}}(e_{i},q_{j})$ be the estimated evaluation score from the LLM and $c_{i}$ be the cost for model $e_{i}$. The ratio is calculated as:
    $\operatorname{ratio}_{ij}=\frac{\operatorname{\hat{eval}}(e_{i},q_{j})}{c_{i}^{\alpha}}$
    where $\alpha$ is a parameter that adjusts the emphasis on cost. A higher $\alpha$ value increases the weight on cost efficiency, thereby favoring cheaper models and preserving more of the budget for subsequent assignments.
  2. 2.
    Sorting and Assignment: For each query, sort the models by their performance-to-cost ratio in descending order. Assign the query to the model with the highest ratio that remains within the available budget.
  3. 3.
    Budget Management: Monitor the accumulated cost. If selecting a model would exceed the budget, move to the next best model in terms of the ratio.
This approach efficiently balances the trade-off between performance and cost, effectively identifying the most suitable model for each query while adhering to budget constraints.
Figure 3: Pipeline of synthetic data generation of Routoo.
Model Accuracy Cost (\$/1M tok) LLaMa2 7B 45.3 0.2 Mistral 7B 64.2 0.2 LLaMa2 13B 54.8 0.26 Mixtral 8x7B 70.6 0.6 LLaMa2 70B 69.9 0.9 Routoo (open-source) 75.87 0.6 GPT3.5 70.0 1.5 GPT4-turbo 86.4 20 Routoo (mix) 84.9 10.2
Table 1: Performance and cost of running LLMs on MMLU benchmark. Accuracy is calculated based on OpenLLM Leaderboard setting Beeching et al. (2023).
3.4 Universe Constructor
Given the vast array of LLMs available for text generation [^4] According to the Huggingface platform (https://huggingface.co), there are nearly 47,000 models available for the text generation task., and considering the practical constraints on the number of models that can be actively served, we propose an optimization approach to construct a complementary universe of models. This universe aims to maximize performance by ensuring that the selected models are complementary to each other.
Figure 4: The performance of different Routoo models and baselines on MMLU benchmark, given different budget limitations.
The objective is to select a subset of LLMs that, collectively, provide the best coverage and performance across a set of queries. Formally, given a set of models $e_{1},e_{2},\ldots,e_{M}$, a set of queries $q_{1},q_{2},\ldots,q_{L}$, and the evaluations of these queries on the models $\operatorname{eval}(e_{i}(q_{j}))$, we seek to find the optimal subset of models that maximizes the average best response for the queries. The optimization problem is formulated as:
$\max_{U\subseteq M,|U|=k}S(U)=\frac{1}{L}\sum_{j=1}^{L}\max_{i\in U}\leavevmode\nobreak\ \operatorname{eval}(e_{i}(q_{j}))$ (4)
where $k$ is the desired number of serving models, $M$ is the set of all available LLMs, and $U$ is the subset of selected models from $M$. The function $S(U)$ represents the highest score achievable by the set $U$, quantifying the combined performance of the selected experts.
Given the submodular nature of the maximum operation in our objective function, a greedy algorithm (see Algorithm 1) can be employed to find an approximate solution with acceptable error bounds Krause and Golovin (2014). This method is particularly effective for large-scale problems where exact optimization is computationally prohibitive.
Further details on the generation of synthetic data for the universe construction are described in Section 4.2.
4 Results and Discussion
4.1 Experiment Setting
Figure 5: Per sub-category performances of our Routoo models and baselines on MMLU benchmark.
Evaluation Setting.
The evaluation of baselines and Routoo models is conducted using MMLU benchmark Hendrycks et al. (2021), which comprises multiple-choice questions across 57 diverse domains, such as mathematics, law, computer science, biology, and US history. To be compatible with OpenLLM Leaderboard Beeching et al. (2023), we use Eleuther AI Harness Gao et al. (2021) to evaluate our models.  [^5] Specifically, the following branch is used: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b0921b636bc36ad05c0b0b0763bd6dd43463. Precisely, it calculates the likelihood of each choice in the question, and selects the answer with maximum likelihood, then ’accuracy’ is used as the evaluation metric. The overall performance is calculated as the average accuracy of the model in 57 domains.
Baselines.
Our comparison includes a range of both open-source and closed-source LLMs. These comprise LLaMa2 Touvron et al. (2023) models with 7b, 13b, and 70b parameters, Mistral 7b Jiang et al. (2023a), Mixtral 8x7b Jiang et al. (2024) (employing token-level MoE), alongside with GPT3.5 and GPT4 OpenAI et al. (2024) as closed-source models.
Architecture Setting.
For the universe constructor, we use all available models of OpenLLM leaderboard Beeching et al. (2023)  [^6] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard. as the underlying LLMs. At the time of writing this paper, there are nearly 2,200 LLMs available on the leaderboard. We set the size of the universe as 56. For the performance predictor, we use Mistral 7b (v0.1) Jiang et al. (2023a)  [^7] https://huggingface.co/mistralai/Mistral-7B-v0.1. as the query encoder. The number of levels in the evaluation score is 2 for MMLU benchmark.  [^8] As accuracy metric is used for MMLU benchmark in OpenLLM leaderboard Beeching et al. (2023). For fine-tuning, we apply LoRA method Hu et al. (2021) with $r=1024,\leavevmode\nobreak\ \alpha=16,\leavevmode\nobreak\ dropout=0.05$ on query, key, and value matrices.
4.2 Training Data Preparation
The data generation pipeline for Routoo is depicted in Figure 3, featuring a query generator, a response generator, and an evaluation system (specifically the MMLU evaluation system). The process begins with the specialized question generator; questions were originated by two approaches: Filtering available open-source datasets and generating queries using GPT4 model OpenAI et al. (2024).
Filtering Available Datasets
: We began by collecting various multiple-choice QA datasets, such as ARC Clark et al. (2018), MC-TEST Richardson et al. (2013), OBQA Mihaylov et al. (2018), RACE Lai et al. (2017), and TruthfulQA Lin et al. (2022). To identify questions with a high training signal (i.e., difficulty), we employed SOLAR-10.7B-v1.0 Kim et al. (2024)  [^9] Available in Huggingface platform: https://huggingface.co/upstage/SOLAR-10.7B-v1.0. We chose this model, as it is performing relatively well on MMLU section of OpenLLM benchmark Beeching et al. (2023). on this aggregated dataset, and calculate the evaluation score. Questions where the model performed poorly were retained, yielding 45,645 challenging training samples from an initial pool of 101,434. Additionally, 10,000 simpler questions were randomly selected from the initial dataset, bringing the total to 55,645 queries.
Generation with GPT4
: To diversify our training dataset further, we generated 20,000 synthetic queries using the GPT4 model. These queries were based on domain and seed data consisting of five random questions from our dataset mentioned above. Detailed specifications of the input prompts used for generation are documented in Appendix B.
After generating the queries, they are passed to the response generator to generate outputs from LLMs in the universe set. Finally, for a given query $q_{j}$ and the response from model $e_{i}$, the evaluation system computes $\operatorname{eval}(e_{i}(q_{j}))$.
The development set of Universe Constructor is created by running a response generator and evaluation system on available LLMs from the OpenLLM benchmark Beeching et al. (2023) (approximately 2,200) for 1,000 randomly selected queries from our training dataset.
4.3 Main Results
Our Routoo Models.
We present two variations of Routoo model: Routoo (open-source) and Routoo (mix). The former refers to a model that leverages 7B, 13B ,and 34B open-source models available on Huggingface as the input to our universe constructor. By additionally integrating GPT4 OpenAI et al. (2024) to underlying models of Routoo (open-source), we create Routoo (mix).
Overall Results.
Comparison of our models and baselines are illustrated in Table 1.101010Inference costs are calculated based on price documentation of the following providers at the time of writing the paper: https://www.together.ai, https://openai.com. For GPT3.5 and GPT4 costs, the average of input and output costs are considered. The cost of Routoo (open-source) is adjusted to match Mixtral 8x7b Jiang et al. (2023a), the best generic open-source LLM (at the time of writing this paper). Our model significantly outperforms Mixtral by 5.27% absolute points with the same inference cost. Also, our underlying models can be executed on 1 GPU (e.g. A100 with 40 GB memory), while Mixtral model requires access to high-end computing resources  (e.g. GPUs with 100 GB memory). Compared to LLaMa2 70B, our model achieves significantly better performance (+6%), while reducing the cost by 33%. Impressively, Routoo (open-source) achieves competitive performance with GPT3.5 while reducing the cost by 73.3%.
Compared to closed-source LLMs, our Routoo (mix) model significantly outperforms GPT3.5 by 14.9% absolute point. Routoo (mix) reaches competitive performance with GPT4, while reducing the cost by almost 50%.
Figure 6: Routing distributions of different variants of Routoo models.
Cost-Aware Decoding.
Given that the cost-aware decoding module can compute different routing distributions within a specified budget, we illustrate the curve of performance-cost in Figure 4 for both open-source and closed-source baselines, alongside with variations of Routoo models. Notably, our Routoo (open-source) reaches the performance of Mixtral 8x7b Jiang et al. (2024) while decreasing the cost by 33%. Interestingly, Routoo (mix) achieves better performance than GPT4 (86.64 vs. 86.40) while reducing the inference cost by nearly 25%.
In summary, as our cost-aware decoder effectively offers an optimal trade-off between cost and performance, both the open-source and mix variants of the Routoo model provide a better performance (or cost) compared to using individual LLMs for queries during inference.
Per Domain Comparison.
To further investigate the source of improvement in our models, we illustrates the distributions of performances on 17 sub-categories Hendrycks et al. (2021) in Figure 5. A standout area of success is in STEM domains, such as mathematics and computer science, where the Routoo particularly excels. This impressive performance is largely attributed to the incorporation of specialized small models (around 7b parameters) that are fine-tuned for tasks in mathematics Yu et al. (2024); Shao et al. (2024) and coding Rozière et al. (2024); Guo et al. (2024) by the community. Furthermore, this approach facilitates the identification of domains where there is a scarcity of experts. Then, future research can focus on improving the performance of these areas by developing domain-specific effective experts.
Routing Distribution.
Figure 6 presents the aggregated percentage usage of expert models by size for each Routoo pricing tier. It reveals that higher-priced options tend to utilize larger models more frequently. A substantial inclusion of effective smaller, 7-billion parameter expert models significantly enhances the cost-to-performance efficiency. This suggests that strategically increasing the use of such smaller experts could offer a more economical solution while maintaining high-quality outputs.
Finally, the cost of training a router is significantly lower than building foundational models e.g. GPT4 OpenAI et al. (2024) and Mixtral Jiang et al. (2024). This could pave a new path for building new frontiers at a much lower cost by integrating knowledge of current LLMs.
5 Conclusion and Future Work
In this paper, we proposed our Routoo architecture, a lightweight LLM-based model that is designed to inteligently routes the input query to the most suitable expert model given other constraints e.g. cost. We evaluated our architecture on MMLU benchmark Hendrycks et al. (2021), which is a MCQA dataset with 57 different domains coverage. Routoo (open-source) achieved competitive performance with Mixtral 8x7b model Jiang et al. (2024) with two-thirds of the inference cost. Increasing the budget limitation allows Routoo (open-source) to outperform Mixtral model by 5% with the same level of cost. By integrating GPT4 OpenAI et al. (2024) model to underlying experts, our Routoo (mix) achieved competitive accuracy with GPT4 while reducing the inference cost by 50%, and even surpassing it while saving 25% of the cost. In general, our Routoo models provide an efficient trade-off between cost and performance during the inference.
In future, the Routoo’s ability to assess and understand the performance of existing models allows researchers to identify gaps in the AI landscape. It pinpoints domains where no existing expert excels or where larger models are inefficiently juggling tasks. This insight is invaluable, enabling us to strategically develop domain-specific models where they are needed most.
Limitations
Due to GPU limitations, we fine-tuned our Routoo models using the LoRA technique Hu et al. (2021). For better performance, full fine-tuning methods could be employed to train our Routoo models. Our cost-aware decoder used a greedy technique to account for inference costs, though more comprehensive methods could better model the relationship between performance and cost. Resource constraints prevented us from using LLMs with 70 billion parameters as our underlying models. Incorporating these larger models, especially in the Routoo (mix) model, yields better performance compared to GPT4. Additionally, integrating other closed-source LLMs e.g. Claude  [^11] https://www.anthropic.com/news/claude-3-family and Gemini  [^12] https://deepmind.google/technologies/gemini/ variants into our Routoo models would increase the diversity of base LLMs.
Appendix A Optimal Routing of Open-source LLMs on MMLU
In this experiment, we utilize all publicly available LLMs from the OpenLLM benchmark Beeching et al. (2023) that have fewer than 34 billion parameters. Given that each query is directed to the most effective and cost-efficient model by an ideal performance predictor module within Routoo, the upper bound for the Routoo model in this scenario is nearly 97.5%. The following table shows the distribution of usage among LLMs of varying sizes:
Model size Usage (%) 7b 66.4 13b 16.1 34b 17.5
Table 2: The Distribution of usage among underlying LLMs with different sizes on MMLU Hendrycks et al. (2021) test set.
The average size of the distribution above could be considered as an abstract model with approximately 13 billion parameters.
Appendix B Synthetic Data Generation by GPT4
The following input prompt is used for generating synthetic data with GPT4 model OpenAI et al. (2024):
[…]
For generation, we used OpenAI API  [^13] https://platform.openai.com/docs/overview. with $max_{tokens}=1000$ and $temperature=1$. SUBJECT is used when the seed samples contain subject field e.g. development set of MMLU benchmark Hendrycks et al. (2021). Seed samples are chosen randomly from datasets introduced in Section 4.2, alongside with development set of MMLU benchmark.

### v3

@@
 Abstract
-Developing foundational large language models (LLMs) is becoming increasingly costly and inefficient.
-Also, closed-source and larger open-source models generally offer better response quality but come with higher inference costs than smaller models.
-In this paper, we introduce Routoo, an architecture designed to optimize the selection of LLMs for specific prompts based on performance, cost, and efficiency.
-Routoo consists of two key components: a performance predictor and a cost-aware decoding.
-The performance predictor is a lightweight LLM that estimates the performance of various underlying LLMs without needing to execute and evaluate them.
-The cost-aware decoding then selects the most suitable model based on these predictions and other constraints like cost and latency.
+LLMs with superior response quality—particularly larger or closed-source models—often come with higher inference costs, making their deployment inefficient and costly.
+Meanwhile, developing foundational LLMs from scratch is becoming increasingly resource-intensive and impractical for many applications.
+To address the challenge of balancing quality and cost, we introduce Routoo, an architecture designed to optimize the selection of LLMs for specific prompts based on performance, cost, and efficiency.
+Routoo provides controllability over the trade-off between inference cost and quality, enabling significant reductions in inference costs for a given quality requirement.
+Routoo comprises two key components: a performance predictor and cost-aware selector.
+The performance predictor is a lightweight LLM that estimates the expected performance of various underlying LLMs on a given prompt without executing them.
+The cost-aware selector module then selects the most suitable model based on these predictions and constraints such as cost and latency, significantly reducing inference costs for the same quality.
 We evaluated Routoo using the MMLU benchmark across 57 domains employing open-source models.
@@
 When integrating GPT4 into our model pool, Routoo nearly matches GPT4’s performance at half the cost and exceeds it with a 25% cost reduction.
-These outcomes highlight Routoo’s potential to create new SOTA in a cost-effective manner by leveraging the collective knowledge of multiple LLMs.
+These outcomes highlight Routoo’s potential to significantly reduce inference costs without compromising quality, and even to establish new state-of-the-art results by leveraging the collective capabilities of multiple LLMs.
 1 Introduction
-Developing foundational models OpenAI et al.
-(2024); Touvron et al.
-(2023); Jiang et al.
-(2023a, 2024) is capital-intensive, necessitating vast computational resources and extensive, high-quality data Minaee et al.
-(2024).
-Furthermore, the field is nearing the upper bounds of network size and data capacity, resulting in progressively marginal enhancements over existing models Udandarao et al.
-(2024).
-This scenario echoes a critical juncture in human advancement, where the ’divide and conquer’ methodology emerges as a viable and scalable alternative.
-This approach entails cultivating domain-specific experts and judiciously harnessing them to forge a composite, high-performance model.
+LLMs have achieved remarkable success across a wide range of natural language processing tasks.
+However, this success comes with a significant trade-off between performance and cost: larger models generally offer better response quality but incur higher inference costs than their smaller counterparts.
+This increased inference cost poses challenges for deploying LLMs in real-world applications where computational resources, latency, and cost-efficiency are critical considerations.
+At the same time, developing foundational LLMs  (OpenAI et al., 2024; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI et al., 2024; Jiang et al., 2024; 2023a; Touvron et al., 2023) from scratch is becoming increasingly capital-intensive, requiring vast computational resources and extensive, high-quality data (Minaee et al., 2024).
+As the field approaches the limits of network size and data capacity, the improvements gained from training ever-larger models are diminishing (Udandarao et al., 2024).
+This situation underscores the need for alternative approaches that can achieve high performance without incurring prohibitive development and inference costs.
+Moreover, many practical tasks do not necessitate the intricate reasoning capabilities of the largest models like GPT-4; instead, they can be efficiently handled by smaller, more cost-effective models (Zaharia et al., 2024; Ding et al., 2024; Shnitzer et al., 2023; Chen et al., 2023).
 The capabilities of existing LLMs appear to be complementary to a significant degree.
-An illustrative case is the MMLU benchmark Hendrycks et al.
-(2021), where selecting the optimal open-source model for each question hypothetically yields an accuracy of 97.5%, at a computational cost akin to a model with 13 billion parameters.
-[^1] Further details are provided in Appendix A.
-In contrast, GPT4 OpenAI et al.
-(2024) achieves an accuracy of 86.4%, while Mixtral 8x7b Jiang et al.
-(2024), as the leading open-source model, reaches 70%.
-These figures suggest considerable scope for integrating LLMs’ knowledge to create new state-of-the-art models.
-Additionally, it is noteworthy that many practical tasks do not require intricate reasoning and can be efficiently addressed by models of moderate complexity; only a minority of tasks demand the advanced capabilities of a model like GPT4 Zaharia et al.
-(2024); Ding et al.
-(2024); Shnitzer et al.
-(2023); Chen et al.
-(2023).
-The landscape of open-source models, with approximately 450,000 entries on Hugging Face, [^2] https://huggingface.co is both dynamic and expansive.
-The rapid proliferation of new models, particularly smaller, domain-focused ones, poses a challenge in tracking and leveraging the latest advancements effectively.
-Figure 1: Routoo architecture has two main components: a performance predictor and a cost-aware decoder.
-The former predicts the accuracy of experts for a specified query, and the latter chooses the underlying LLM and generates the response by considering the cost and efficiency of each model.
-The Universe Constructor builds a complementary set of underlying models.
-We propose Routoo, an architecture to learn to leverage a universe of trained LLMs automatically and dynamically to create a better-performing model.
-The cornerstone of Routoo is a lightweight LLM that predicts the performance of each underlying LLM on queries without the need for actual execution and evaluation.
-Based on these predictions, a cost-aware decoding algorithm selects the most suitable model, prioritizing factors such as inference cost and speed.
-For instance, when faced with a task that is predicted to be performed nearly equally well by a 7 billion parameter model or a more extensive 70 billion parameter model, the Routoo will opt for the former when factors like speed and cost-efficiency are prioritized.
+For example, on the MMLU benchmark (Hendrycks et al., 2021), selecting the optimal open-source model for each question could hypothetically yield an accuracy of 97.5% at a computational cost similar to that of a 13-billion-parameter model (see Appendix  A for details).
+In contrast, GPT-4 (OpenAI et al., 2024) achieves an accuracy of 86.4%, while Mixtral 8×7b (Jiang et al., 2024), one of the leading open-source models, [^1] As of the time of writing this paper.
+reaches 70%.
+These observations suggest that integrating the knowledge of multiple LLMs could lead to new state-of-the-art models without the need to train from scratch while significanly reducing the inference cost.
+However, effectively leveraging the vast and rapidly growing ecosystem of LLMs poses significant challenges.
+With approximately 450,000 models available on Hugging Face, [^2] https://huggingface.co keeping track of the latest advancements and integrating them efficiently is a non-trivial task.
+Traditional approaches like Mixture of Experts (MoE) models (Abdin et al., 2024; Lieber et al., 2024; Jiang et al., 2024; Fedus et al., 2021; Shazeer et al., 2017) are limited by the need to load all expert parameters onto a single high-end machine, hindering scalability and flexibility.
+There is a need for an architecture that can dynamically and efficiently leverage multiple existing LLMs to optimize performance while controlling inference costs.
+To address these challenges, we propose Routoo , an architecture designed to optimize the selection of LLMs for specific prompts based on performance, cost, and efficiency.
+Routoo provides controllability over the trade-off between inference cost and quality, enabling significant reductions in inference costs for a given quality requirement.
+By intelligently leveraging a universe of trained LLMs, Routoo addresses both the inference cost problem and the development cost of building LLMs from scratch, creating a composite high-performance model without the need for extensive retraining.
+Routoo comprises two key components: a performance predictor and cost-aware selector.
+The performance predictor is a lightweight LLM that estimates the expected performance of various underlying LLMs on a given query without executing them.
+Based on these predictions and constraints such as cost and latency, the cost-aware selector module selects the most suitable model.
+For instance, when a task is predicted to be performed nearly equally well by a smaller model or a larger, more expensive model, Routoo will opt for the smaller model when speed and cost-efficiency are prioritized.
 This approach ensures optimal resource utilization without compromising on quality.
-Our architecture marks a significant departure from traditional Mixture of Experts (MoE) models Lieber et al.
-(2024); Jiang et al.
-(2024); Fedus et al.
-(2021); Shazeer et al.
-(2017).
+Our architecture marks a significant departure from traditional MoE.
 While MoE relies on gating over various expert sub-networks within each layer to predict the next token, it requires all expert parameters to be loaded onto a single, high-end machine.
@@
 In contrast, each ’expert’ within our system operates independently and can be hosted on different machines, potentially utilizing a different neural network architecture.
-This flexibility means we can incorporate a vast array of experts, ranging from those specializing in system-level Java programming to those adept at curating travel experiences in London.
-Moreover, Routoo can be tuned to leverage only smaller models, which is particularly beneficial for users without access to high-end computing resources (e.g., GPUs with 100GB).
-Furthermore, our architecture facilitates easy integration of additional optimization criteria, including cost, speed, and privacy considerations.
-We evaluate our Routoo on MMLU benchmark Hendrycks et al.
-(2021).
-We show that it achieves competitive performance with Mixtral 8x7b Jiang et al.
-(2024) while only consuming two-thirds of its inference cost.
+This flexibility enables Routoo to incorporate a vast array of experts, ranging from specialized domain models to general-purpose ones, and to scale to a large number of experts without the limitations imposed by MoE models.
+We evaluate our Routoo on MMLU benchmark (Hendrycks et al., 2021).
+We show that it achieves competitive performance with Mixtral 8x7b (Jiang et al., 2024) while only consuming two-thirds of its inference cost.
 Increasing the cost budget allows Routoo to outperform Mixtral by 5% at the same cost level, reaching an accuracy of 75.9%.
-By adding GPT4 OpenAI et al.
-(2024) as one of the underlying experts, our Routoo achieves competitive performance with GPT4 at half the cost and exceeds it with a 25% cost reduction.
+By adding GPT4 (OpenAI et al., 2024) as one of the underlying experts, our Routoo achieves competitive performance with GPT4 at half the cost and exceeds it with a 25% cost reduction.
+These outcomes highlight Routoo’s potential to significantly reduce inference costs without compromising quality, and even to establish new state-of-the-art results by leveraging the collective capabilities of multiple LLMs.
+By providing controllability over cost and quality trade-offs, Routoo offers an efficient approach to high-performance language modeling without the need for expensive model training from scratch.
 To summarize, our contributions are:
-  • We propose Routoo, an LLM-based system designed to intelligently identify the best-performing LLM for a given query while considering constraints (e.g.
-cost), effectively integrating knowledge of multiple LLMs to create a new state-of-the-art model.
-  • We evaluate our architecture on MMLU benchmark, and significantly outperform Mixtral 7x8b by 5% with similar inference cost.
+  • We propose Routoo, an LLM-based system that intelligently identifies the best-performing LLM for a given query while considering constraints such as cost and latency, effectively integrating the knowledge of multiple LLMs to create a high-performance model without the need for training from scratch.
+  • We evaluate our architecture on MMLU benchmark, and significantly outperform Mixtral 8x7b by 5% with similar inference cost.
 Also, our Routoo achieves competitive performance with GPT4 at half the inference cost and surpasses it by reducing the inference cost by 25%.
 2 Related Work
-LLM Benchmarks.
-In recent years, many benchmarks Wang et al.
-(2024); Liang et al.
-(2023); Beeching et al.
-(2023); Gao et al.
-(2021); Hendrycks et al.
-(2021) are created to identify the capability of LLMs.
-The Massive Multitask Language Understanding  (MMLU) benchmark Hendrycks et al.
-(2021) contains multiple-choice question-answer (MCQA) pairs from various branches of knowledge.
-This includes 57 domains, including mathematics, US history, computer science, and law.
-OpenLLM leaderboard Beeching et al.
-(2023) includes several question-answering datasets, e.g.
-HellaSwag Zellers et al.
-(2019), Truthfulqa Lin et al.
-(2022), Winogrande Sakaguchi et al.
-(2019), GSM8k Cobbe et al.
-(2021), ARC Clark et al.
-(2018) and MMLU, which is used to compare open-source LLMs.
-HELM Liang et al.
-(2023) consists of 42 scenarios covering a variety of use cases.
-In this work, we utilized the MMLU benchmark Hendrycks et al.
-(2021) to assess our models.
-This benchmark spans several domains and is recognized as one of the challenging benchmarks on the OpenLLM leaderboard Lu et al.
-(2023).
+Mixture-of-Experts.
+MoE architecture includes a gating mechanism to integrate various expert sub-networks within each layer, guiding the prediction of the next token (Shazeer et al., 2017).
+This approach is used in recent foundational models, including Mixtral 8x7b (Jiang et al., 2024).
+Sukhbaatar et al.
+(2024) fine-tuned LLaMa 7b (Touvron et al., 2023) on four different domains to create domain-specific expert LLMs.
+They then proposed combining the FFNs of these expert LLMs to construct an MoE.
+Tang et al.
+(2024) proposed merging most parameters while scaling up MLP layrers of Transformer into a weight-ensembling MoE module that dynamically integrates shared and task-specific knowledge based on the input.
+Wang et al.
+(2024) introduces a fusion gate at each Transformer layer to generate weights for a weighted average of outputs from a set of pre-trained LoRAs.
+Figure 1: Routoo architecture has three main components: a performance predictor, cost-aware selector, and universe constructor.
+The universe constructor identifies the optimal set of complementary models from all available models.
+The performance predictor predicts the correctness of experts for a specified query, and the cost-aware selector chooses the underlying model by considering the cost and efficiency of each model.
 Model Selection.
 Previous works in selecting the best LLMs mainly focus on identifying the one that generates the most optimal output for a given input.
-Liu and Liu (2021); Ravaut et al.
+Liu & Liu (2021); Ravaut et al.
 (2022) proposed specialized scoring or re-ranking models that can be used for the generation tasks (summarisation, here).
-Speculative decoding Kim et al.
-(2023); Leviathan et al.
-(2023) accelerates the decoding of expensive models by using small, efficient decoders for the ’easy’ steps.
+Speculative decoding (Kim et al., 2023; Leviathan et al., 2023) accelerates the decoding of expensive models by using small, efficient decoders for the ’easy’ steps.
 This approach can complement routing methods.
-LLM-BLENDER Jiang et al.
-(2023b) is an ensembling framework to reach better performance by mixing the results of LLMs with a ranking module, followed by a generation module to generate from top candidates of the ranker.
-FrugalGPT Chen et al.
-(2023) executes experts sequentially until an expert reaches the acceptable generation performance.
-HybridLLM Ding et al.
-(2024) proposes a routing approach to direct queries to the suitable expert, utilizing two underlying LLMs.
+LLM-BLENDER (Jiang et al., 2023b) is an ensembling framework to reach better performance by mixing the results of LLMs with a ranking module, followed by a generation module to generate from top candidates of the ranker.
+FrugalGPT (Chen et al., 2023) executes experts sequentially until an expert reaches the acceptable generation performance.
+ZOOTER (Lu et al., 2024) proposed a reward-guided routing approach that distills rewards from training queries to train a routing function.
+HybridLLM (Ding et al., 2024) and RouteLLM (Ong et al., 2024) propose a routing approach to direct queries to the suitable expert, utilizing one strong and one weak LLMs.
 Different from previous work, our Routoo identifies the most suitable expert without executing the inference of underlying LLMs.
-Additionally, our approach can be efficiently generalized to scenarios involving many underlying LLMs.
+Additionally, our approach can be efficiently generalized to scenarios involving many underlying LLMs (not necessarily a tuple of weak and strong LLMs).
 3 Architecture
 3.1 Problem Formulation
-Given a set of $N$ LLM $e_{1},e_{2},\ldots,e_{N}$ and a set of $m$ queries $q_{1},q_{2},\ldots,q_{m}$, Our goal is to assign the most cost-effective model $e_{i}$ to each query $q_{j}$that can accurately answer the query.
-For each model-query pair, we define an evaluation function $\operatorname{eval}(e_{i}(q_{j}))$ [^3] Evaluation function can have the reference as an optional input., which returns a score from 0 to $K-1$ indicating the correctness of the model’s response, where 0 is unacceptable and $K-1$ is optimal.
-Each response also incurs a cost $\operatorname{cost}(e_{i}(q_{j}))$.
+Given a set of $M$ models $m_{1},m_{2},\ldots,m_{M}$ and a set of $N$ queries $q_{1},q_{2},\ldots,q_{N}$, [^3] Section 3.4 will provide a detailed explanation of how we build the set of models from the pool of all available models.
+our goal is to assign the most cost-effective model $m_{i}$ to each query $q_{j}$ that can accurately answer the query.
+Correctness and cost scores of each query-model pair are calculated as:
+$s_{m_{i},q_{j}}=\operatorname{Eval}(m_{i}(q_{j}))$
+$c_{m_{i},q_{j}}=\operatorname{Cost}(m_{i},q_{j})$
+where $\operatorname{Eval}(.)$ computes the correctness score given the generated response $m_{i}(q_{j})$ of model $m_{i}$ for query $q_{j}$.
+$s_{m_{i},q_{j}}$ is an integer that ranges from 0 to $K-1$, indicating the correctness of the model’s response, where 0 is unacceptable, and $K-1$ is optimal.
+The cost of executing model $m_{i}$ for query $q_{j}$ is calculated by $\operatorname{Cost}(.)$ function.
 The objective is to maximize the total correctness scores across all queries while keeping the total cost within a budget $B$:
-$\max_{\pi}$ $\ \sum_{j=1}^{m}\operatorname{eval}(\pi(q_{j})(q_{j}))$
-s.t. $\ \sum_{j=1}^{m}\operatorname{cost}(\pi(q_{j})(q_{j}))\leq B$
-Here, $\pi$ is the function that assigns each query to a model, based on both cost and correctness.
+$\max_{\pi}$ $\ \sum_{j=1}^{N}s_{\pi(q_{j}),q_{j}}$
+s.t. $\ \sum_{j=1}^{N}c_{\pi(q_{j}),q_{j}}\leq B$
+Here, $\pi(.)$ is the function that assigns each query to a model from the set of $M$ models, while considering both cost and correctness.
 In this formulation, the cost can be multi-faceted, encompassing aspects like computational cost, speed, etc.
@@
   1. 1.
-    Performance Prediction: This component approximates the evaluation score $\operatorname{eval}(e_{i}(q_{j}))$ for each model on each query without actual execution.
+    Performance Predictor.
+This component approximates the correctness score $s_{m_{i},q_{j}}$ for each model on each query.
+Given a model $m_{i}$ and a query $q_{j}$, it estimates the model’s performance without generating the response for the query, thereby significantly reducing the inference cost.
   2. 2.
-    Cost-Aware Decoder: This step selects the best model for each query, balancing accuracy and cost-effectiveness using the performance predictions from the previous component.
-These components work together to dynamically and efficiently allocate models to queries, as illustrated in Figure 1. In the following sections, we will delve deeper into each component.
+    Cost-Aware Selector. This module selects the best model for each query, balancing accuracy and cost-effectiveness by using correctness estimations of performance predictor.
+These two components work together to dynamically and efficiently allocate models to queries.
+Additionally, given the vast array of LLMs available for text generation, [^4] According to the Huggingface platform (https://huggingface.co), there are nearly 47,000 models available for the text generation task.
+and considering the practical constraints on the number of models that can be actively served, we introduce Universe Constructor, which builds an initial complementary universe of models ($m_{1},m_{2},\ldots,m_{M}$) from a set of all available models.
+This universe aims to maximize performance by ensuring that the selected models are complementary to each other.
+All three components are illustrated in Figure 1. In the following sections, we will delve deeper into each component.
 3.2 Performance Predictor
-Figure 2: The performance predictor: Given the query and an LLM, it outputs the performance prediction $\operatorname{\hat{eval}}(e_{i},q_{j})$.
 The performance predictor is a lightweight LLM designed to estimate the effectiveness of each underlying LLM for a given query.
-It predicts the evaluation score, representing the expected response quality of a model to a query.
+It estimates the output of evaluation function ($s_{m_{i},q_{j}}$, introduced in Section 3.1) without executing model $m_{i}$ on query $q_{j}$.
 This prediction is formulated as:
-$\operatorname{\hat{eval}}(e_{i},q_{j})=\operatorname{Pred}(e_{i},q_{j}),$ (1)
-where $\operatorname{\hat{eval}}(e_{i},q_{j})$ denotes the predicted score of model $e_{i}$ for query $q_{j}$, and $\operatorname{Pred}(e_{i},q_{j})$ is the predictive model.
-The process, illustrated in Figure 2, involves:
-$\begin{split}\begin{cases}h_{q_{j}}=\operatorname{Enc}(q_{j})\\ h_{e_{i}}=\operatorname{Emb}(e_{i})\\ s_{q_{j},e_{i}}=\operatorname{Linear}(h_{q_{j}}-h_{e_{i}})\end{cases}\end{split}$
-where $\operatorname{Enc}(.)$ is the encoder of the input query.
-Inspired by Radford et al.
-(2019), we use a decoder-only model as the encoder and extract the embedding of the last token as the representation of the input query.
-$Emb(.)$ is an embedding layer, where each embedding is assigned to a specific model.
-$Linear(.)$ is a linear layer to convert the combination of query and model embeddings to the logits of the evaluation score.
-$s_{q_{i},e_{j}}$ is a list of predicted evaluation scores for a given model and query.
-The size of this list depends on the number of levels (K) in the evaluation score e.g.
-2 for binary evaluation scores.
-The model is trained by minimising the cross-entropy loss function Good (1952) of $s_{q_{i},e_{j}}$ and $\operatorname{eval}(e_{j}(q_{i}))$ over all $N$ expert models, and $m$ queries as:
-$\operatorname{min}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{N}\operatorname{CE}(s_{q_{i},e_{j}},\operatorname{eval}(e_{j}(q_{i})))$ (2)
-where $\operatorname{CE}(.)$ is the cross-entropy loss.
-During the inference time, $\operatorname{argmax}(.)$ is applied for a given model and query as:
-$\operatorname{\hat{eval}}(e_{i},q_{j})=\operatorname{argmax}_{(0,...,K-1)}(s_{q_{j},e_{i}})$ (3)
-where $\operatorname{argmax}(.)$ is applied over the levels of the evaluation score.
-3.3 Cost-Aware Decoder
-The second phase of our architecture involves the decoding step, where the estimated scores from the performance predictor are used to determine the optimal assignment of models to queries.
-Our objective is to maximize the overall effectiveness of the responses within a given budget constraint.
-The optimization problem is formulated as follows:
-$\max_{\pi}$ $\ \sum_{j=1}^{m}\operatorname{\hat{eval}}(\pi(q_{j}),q_{j})$
-s.t. $\ \sum_{j=1}^{m}\operatorname{cost}(\pi(q_{j})(q_{j}))\leq B$
-where $\pi(.)$ determines the expert assignments and $B$ represents the predefined budget constraint.
-We propose a greedy algorithm to approximate this optimization.
-Noting the challenges of accurately estimating the exact cost of a model for a specific query, instead, we consider the average cost $c_{i}$ for a model $e_{i}$ responding to an average length input and response.
+$\hat{s}_{m_{i},q_{j}}=\operatorname{Pred}(m_{i},q_{j})$
 Result: Find the set U of size k that maximizes $S(U)$ according to equation (2)
-Initialize an empty set $U=\{\}$
+Initialize an empty set $U=\{\}$, and set max budget to M
 Set budget to 0
-while budget < k do
-       1. Identify the model $e_{i}^{*}$ that, when added to $U$ forming $U^{*}$, maximizes $S(U^{*})$
-       2. Update $U$ to $U^{*}$
+while budget $<$ M do
+       1. Identify the model $m_{i}^{*}$ that when added to $U$ forming $U^{*}$ which maximizes $S(U^{*})$
+       2. Update $U$ by adding $m_{i}^{*}$
        3. Increment budget by 1
@@
 end while
-Algorithm 1 Universe Construction.
+Algorithm 1 Greedy algorithm for the optimization of universe construction.
+where $\operatorname{Pred}(m_{i},q_{j})$ is the predictive model.
+Similar to $s_{m_{i},q_{j}}$, $\hat{s}_{m_{i},q_{j}}$ is an integer that ranges from 0 to $K-1$, 0 indicates an unacceptable result and $K-1$ represents the optimal outcome.
+The process of $\operatorname{Pred}(m_{i},q_{j})$ is as follows:
+$\begin{split}\begin{cases}h_{q_{j}}=\operatorname{Enc}(q_{j})\\ h_{m_{i}}=\operatorname{Emb}(m_{i})\\ \hat{s}_{m_{i},q_{j}}=\operatorname{Linear}(h_{q_{j}}-h_{m_{i}})\end{cases}\end{split}$
+where $\operatorname{Enc}(.)$ is the encoder of the input query.
+Inspired by Radford et al.
+(2019), we use a decoder-only LLM as the query encoder and extract the embedding of the last token as the representation of the input query ($h_{q_{j}}\in\mathbb{R}^{h}$).
+$\operatorname{Emb}(.)$ is an embedding layer, where each embedding is assigned to a specific model.
+As a result, $h_{e_{i}}\in\mathbb{R}^{h}$ is the embedding representation of model $e_{i}$.
+$h$ is the hidden representation of the decoder-only model used.
+$\operatorname{Linear}(.)$ function is a projection matrix ($\mathbb{R}^{h\times K}$), computing the estimation score $\hat{s}_{m_{i},q_{j}}$.
+The model is trained by minimising the cross-entropy loss function (Good, 1952) of $s_{m_{i},q_{j}}$ and $\hat{s}_{m_{i},q_{j}}$ over all $M$ models, and $N$ queries as:
+$\operatorname{min}\frac{1}{N\times M}\sum_{i=1}^{M}\sum_{j=1}^{N}\operatorname{CE}(s_{m_{i},q_{j}},\hat{s}_{m_{i},q_{j}})$ (1)
+where $\operatorname{CE}(.)$ is the cross-entropy loss.
+During the inference time, $\operatorname{argmax}(.)$ function is applied for a given model and query.
+3.3 Cost-Aware Selector
+The second phase of our architecture involves the selection step, where the estimated scores from the performance predictor are used to determine the optimal assignment of models to queries.
+Our objective is to maximize the overall effectiveness of the responses within a given budget constraint.
+The optimization problem is formulated as follows:
+$\max_{\pi}$ $\ \sum_{j=1}^{N}\hat{s}_{\pi(q_{j}),q_{j}}$
+s.t. $\ \sum_{j=1}^{N}c_{\pi(q_{j}),q_{j}}\leq B$
+where $\pi(.)$ determines the model assignments and $B$ represents the predefined budget constraint.
+We propose a greedy algorithm to approximate this optimization.
+Noting the challenges of accurately estimating the exact cost of a model for a specific query, instead, we consider the average cost $c_{i}$ for a model $m_{i}$ responding to an average length query and response.
 The algorithm includes the following steps:
   1. 1.
-    Performance-to-Cost Ratio: For each query $q_{j}$ and model $e_{i}$, calculate the performance-to-cost ratio.
-Let $\operatorname{\hat{eval}}(e_{i},q_{j})$ be the estimated evaluation score from the LLM and $c_{i}$ be the cost for model $e_{i}$.
+    Performance-to-Cost Ratio: For each query $q_{j}$ and model $m_{i}$, calculate the performance-to-cost ratio.
+Let $\hat{s}_{m_{i},q_{j}}$ be the estimated correctness score of model $m_{i}$ for query $q_{j}$, and $c_{i}$ be the cost for model $m_{i}$.
 The ratio is calculated as:
-    $\operatorname{ratio}_{ij}=\frac{\operatorname{\hat{eval}}(e_{i},q_{j})}{c_{i}^{\alpha}}$
+    $r_{m_{i},q_{j}}=\frac{\hat{s}_{m_{i},q_{j}}}{c_{i}^{\alpha}}$
     where $\alpha$ is a parameter that adjusts the emphasis on cost.
@@
 This approach efficiently balances the trade-off between performance and cost, effectively identifying the most suitable model for each query while adhering to budget constraints.
-Figure 3: Pipeline of synthetic data generation of Routoo.
 Model Accuracy Cost (\$/1M tok) LLaMa2 7b 45.3 0.2 Mistral 7b 64.2 0.2 LLaMa2 13b 54.8 0.26 Mixtral 8x7b 70.6 0.6 LLaMa2 70b 69.9 0.9 Routoo (open-source) 75.87 0.6 GPT3.5 70.0 1.5 GPT4-turbo 86.4 20 Routoo (mix) 84.9 10.2
@@
 3.4 Universe Constructor
-Given the vast array of LLMs available for text generation [^4] According to the Huggingface platform (https://huggingface.co), there are nearly 47,000 models available for the text generation task., and considering the practical constraints on the number of models that can be actively served, we propose an optimization approach to construct a complementary universe of models.
-This universe aims to maximize performance by ensuring that the selected models are complementary to each other.
-Figure 4: The performance of different Routoo models and baselines on MMLU benchmark, given different budget limitations.
-The objective is to select a subset of LLMs that, collectively, provide the best coverage and performance across a set of queries.
-Formally, given a set of models $e_{1},e_{2},\ldots,e_{M}$, a set of queries $q_{1},q_{2},\ldots,q_{L}$, and the evaluations of these queries on the models $\operatorname{eval}(e_{i}(q_{j}))$, we seek to find the optimal subset of models that maximizes the average best response for the queries.
-The optimization problem is formulated as:
-$\max_{U\subseteq M,|U|=k}S(U)=\frac{1}{L}\sum_{j=1}^{L}\max_{i\in U}\leavevmode\nobreak\ \operatorname{eval}(e_{i}(q_{j}))$ (4)
-where $k$ is the desired number of serving models, $M$ is the set of all available LLMs, and $U$ is the subset of selected models from $M$.
-The function $S(U)$ represents the highest score achievable by the set $U$, quantifying the combined performance of the selected experts.
-Given the submodular nature of the maximum operation in our objective function, a greedy algorithm (see Algorithm 1) can be employed to find an approximate solution with acceptable error bounds Krause and Golovin (2014).
+With the abundance of models and the practical limitations on hosting several models, we introduce an optimization approach to build a complementary subset of models ($m_{1},m_{2},\ldots,m_{M}$).
+The goal is to select models that are complementary to achieve the highest performance, given a predefined limitation on number of serving models.
+The objective is to select a subset of models that collectively provide the best coverage and performance across a set of queries.
+Formally, given a big pool of models $\Omega:(m_{1},m_{2},\ldots,m_{\hat{M}})$, a set of queries $Q:(q_{1},q_{2},\ldots,q_{L})$, and correctness scores of the models for these queries $s_{m_{i},q_{j}}$, we seek to find the optimal subset of models, that maximizes the following optimization problem:
+$\max_{U\subseteq\Omega,|U|=M}S(U)=\frac{1}{L}\sum_{j=1}^{L}\max_{i\in U}~{}s_{m_{i},q_{j}}$ (2)
+where $M$ is the desired number of serving models, and $U$ is a subset of selected models from $\Omega$.
+The function $S(U)$ represents the highest score achievable by the set $U$, quantifying the combined performance of the selected models.
+The optimal subset $S(U^{*})$ is used as the set of models ($m_{1},m_{2},\ldots,m_{M})$ in the optimization of Section 3.1.
+Given the submodular nature of the maximum operation in our objective function (Equation 2), a greedy algorithm (as illustrated in Algorithm 1) can be employed to find an approximate solution with acceptable error bounds (Krause & Golovin, 2014).
 This method is particularly effective for large-scale problems where exact optimization is computationally prohibitive.
-Further details on the generation of synthetic data for the universe construction are described in Section 4.2.
 4 Results and Discussion
 4.1 Experiment Setting
-Figure 5: Per sub-category performances of our Routoo models and baselines on MMLU benchmark.
 Evaluation Setting.
-The evaluation of baselines and Routoo models is conducted using MMLU benchmark Hendrycks et al.
-(2021), which comprises multiple-choice questions across 57 diverse domains, such as mathematics, law, computer science, biology, and US history.
-To be compatible with OpenLLM Leaderboard Beeching et al.
-(2023), we use Eleuther AI Harness Gao et al.
-(2021) to evaluate our models.
+The evaluation of baselines and Routoo models is conducted using MMLU benchmark (Hendrycks et al., 2021), which comprises multiple-choice questions across 57 diverse domains, such as mathematics, law, computer science, biology, and US history.
+To be compatible with OpenLLM Leaderboard (Beeching et al., 2023), we use Eleuther AI Harness (Gao et al., 2021) to evaluate our models.
 [^5] Specifically, the following branch is used: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b0921b636bc36ad05c0b0b0763bd6dd43463.
@@
 The overall performance is calculated as the average accuracy of the model in 57 domains.
+Routoo models will be made publicly available for the AI community to test on various benchmarks.
+Figure 2: The performance of different Routoo models and baselines on MMLU benchmark, given different budget limitations.
 Baselines.
 Our comparison includes a range of both open-source and closed-source LLMs.
-These comprise LLaMa2 Touvron et al.
-(2023) models with 7b, 13b, and 70b parameters, Mistral 7b Jiang et al.
-(2023a), Mixtral 8x7b Jiang et al.
-(2024) (employing token-level MoE), alongside with GPT3.5 and GPT4 OpenAI et al.
-(2024) as closed-source models.
+These comprise LLaMa2 (Touvron et al., 2023) models with 7b, 13b, and 70b parameters, Mistral 7b (Jiang et al., 2023a), Mixtral 8x7b (Jiang et al., 2024) (employing token-level MoE), alongside with GPT3.5 and GPT4 (OpenAI et al., 2024) as closed-source models.
+We could not find any publicly available trained models for routing or model integration.
+To the best of our knowledge, there are currently no open-source models that perform routing or dynamic selection among multiple LLMs that we could directly compare against.
 Architecture Setting.
-For the universe constructor, we use all available models of OpenLLM leaderboard Beeching et al.
-(2023)  [^6] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard.
-as the underlying LLMs.
-At the time of writing this paper, there are nearly 2,200 LLMs available on the leaderboard.
-We set the size of the universe as 56.
-For the performance predictor, we use Mistral 7b (v0.1) Jiang et al.
-(2023a)  [^7] https://huggingface.co/mistralai/Mistral-7B-v0.1.
+For the pool of universe constructor ($\Omega$), we use top 1,000 available models of OpenLLM leaderboard (Beeching et al., 2023)  [^6] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard., Then, we set the maximum number of serving models ($M$) to 56 in Equation 2, meaning that we use 56 models for the routing optimization, defined in Section 3.1.
+For the performance predictor, we use Mistral 7b (v0.1) (Jiang et al., 2023a)  [^7] https://huggingface.co/mistralai/Mistral-7B-v0.1.
 as the query encoder.
-The number of levels in the evaluation score is 2 for MMLU benchmark.
-[^8] As accuracy metric is used for MMLU benchmark in OpenLLM leaderboard Beeching et al.
-(2023).
-For fine-tuning, we apply LoRA method Hu et al.
-(2021) with $r=1024,\leavevmode\nobreak\ \alpha=16,\leavevmode\nobreak\ dropout=0.05$ on query, key, and value matrices.
+The number of levels in the evaluation score ($K$) is set to 2 for MMLU benchmark, meaning that $s_{m_{i},q_{j}}$ and $\hat{s}_{m_{i},q_{j}}$ are either 0 or 1.
+[^8] As accuracy metric is used for MMLU benchmark in OpenLLM leaderboard (Beeching et al., 2023).
+For fine-tuning, we apply LoRA method (Hu et al., 2021) with $r=1024,~{}\alpha=16,~{}\text{dropout}=0.05$ on query, key, and value matrices.
 4.2 Training Data Preparation
-The data generation pipeline for Routoo is depicted in Figure 3, featuring a query generator, a response generator, and an evaluation system (specifically the MMLU evaluation system).
-The process begins with the specialized question generator; questions were originated by two approaches: Filtering available open-source datasets and generating queries using GPT4 model OpenAI et al.
-(2024).
-Filtering Available Datasets
-: We began by collecting various multiple-choice QA datasets, such as ARC Clark et al.
-(2018), MC-TEST Richardson et al.
-(2013), OBQA Mihaylov et al.
-(2018), RACE Lai et al.
-(2017), and TruthfulQA Lin et al.
-(2022).
-To identify questions with a high training signal (i.e., difficulty), we employed SOLAR-10.7B-v1.0 Kim et al.
-(2024)  [^9] Available in Huggingface platform: https://huggingface.co/upstage/SOLAR-10.7B-v1.0.
-We chose this model, as it is performing relatively well on MMLU section of OpenLLM benchmark Beeching et al.
-(2023).
+Since the MMLU dataset (Hendrycks et al., 2021) lacks a training set, we build synthetic data to train the models.
+The synthetic questions are originated by two approaches: Filtering available open-source datasets and generating synthetic queries using GPT4 model (OpenAI et al., 2024).
+Filtering Available Datasets.
+We begin by collecting various multiple-choice QA datasets, such as ARC (Clark et al., 2018), MC-TEST (Richardson et al., 2013), OBQA (Mihaylov et al., 2018), RACE (Lai et al., 2017), and TruthfulQA (Lin et al., 2022).
+To identify questions with a high training signal (i.e., difficulty), we employed SOLAR-10.7B-v1.0 (Kim et al., 2024)  [^9] Available in Huggingface platform: https://huggingface.co/upstage/SOLAR-10.7B-v1.0.
+We chose this model, as it is performing relatively well on MMLU section of OpenLLM benchmark (Beeching et al., 2023).
 on this aggregated dataset, and calculate the evaluation score.
-Questions where the model performed poorly were retained, yielding 45,645 challenging training samples from an initial pool of 101,434.
+[^10] MMLU evaluation system (introduced in Section 4.1) is used.
+Queries on which the model under-performed (accuracy of 0) were retained, resulting in a set of 45,645 challenging training samples from an initial pool of 101,434.
 Additionally, 10,000 simpler questions were randomly selected from the initial dataset, bringing the total to 55,645 queries.
-Generation with GPT4
-: To diversify our training dataset further, we generated 20,000 synthetic queries using the GPT4 model.
-These queries were based on domain and seed data consisting of five random questions from our dataset mentioned above.
+Figure 3: Per sub-category performances of our Routoo models and baselines on MMLU benchmark.
+Generation with GPT4.
+To diversify our training dataset further, we generated 20,000 synthetic queries using GPT4 model.
+These queries are generated by using seed data, which randomly sampled from the dataset mentioned above.
 Detailed specifications of the input prompts used for generation are documented in Appendix B.
-After generating the queries, they are passed to the response generator to generate outputs from LLMs in the universe set.
-Finally, for a given query $q_{j}$ and the response from model $e_{i}$, the evaluation system computes $\operatorname{eval}(e_{i}(q_{j}))$.
-The development set of Universe Constructor is created by running a response generator and evaluation system on available LLMs from the OpenLLM benchmark Beeching et al.
-(2023) (approximately 2,200) for 1,000 randomly selected queries from our training dataset.
+Final set of queries contains nearly 75,000 samples.
+Datatrove library [^11] https://github.com/huggingface/datatrove.
+is used for decontamination.
+The set of queries in Universe Constructor ($Q:(q_{1},q_{2},\ldots,q_{L})$) is created by randomly sampling 1,000 queries from the training dataset.
 4.3 Main Results
@@
 The former refers to a model that leverages 7b, 13b ,and 34b open-source models available on Huggingface as the input to our universe constructor.
-By additionally integrating GPT4 OpenAI et al.
-(2024) to underlying models of Routoo (open-source), we create Routoo (mix).
+By additionally integrating GPT4 (OpenAI et al., 2024) to underlying models of Routoo (open-source), we create Routoo (mix).
 Overall Results.
@@
 For GPT3.5 and GPT4 costs, the average of input and output costs are considered.
-The cost of Routoo (open-source) is adjusted to match Mixtral 8x7b Jiang et al.
-(2023a), the best generic open-source LLM (at the time of writing this paper).
-Our model significantly outperforms Mixtral by 5.27% absolute points with the same inference cost.
+The cost of Routoo (open-source) is adjusted to match Mixtral 8x7b (Jiang et al., 2023a), the best generic open-source LLM (at the time of writing this paper).
+Our model significantly outperforms Mixtral 8x7b by 5.27% absolute points with the same inference cost.
 Also, our underlying models can be executed on 1 GPU (e.g.
@@
 Routoo (mix) reaches competitive performance with GPT4, while reducing the cost by almost 50%.
-Figure 6: Routing distributions of different variants of Routoo models.
-Cost-Aware Decoding.
-Given that the cost-aware decoding module can compute different routing distributions within a specified budget, we illustrate the curve of performance-cost in Figure 4 for both open-source and closed-source baselines, alongside with variations of Routoo models.
-Notably, our Routoo (open-source) reaches the performance of Mixtral 8x7b Jiang et al.
-(2024) while decreasing the cost by 33%.
-Interestingly, Routoo (mix) achieves better performance than GPT4 (86.64 vs.
-86.40) while reducing the inference cost by nearly 25%.
-In summary, as our cost-aware decoder effectively offers an optimal trade-off between cost and performance, both the open-source and mix variants of the Routoo model provide a better performance (or cost) compared to using individual LLMs for queries during inference.
+Figure 4: Routing distributions of different variants of Routoo models. The inference cost is provided for 1 million tokens.
+Cost-Aware Selection.
+[…]
+In summary, our cost-aware selector effectively offers an optimal trade-off between cost and performance, enabling both open-source and mix variants of the Routoo model to outperform individual LLMs in terms of performance (or cost) during inference.
 Per Domain Comparison.
-To further investigate the source of improvement in our models, we illustrates the distributions of performances on 17 sub-categories Hendrycks et al.
-(2021) in Figure 5.
-A standout area of success is in STEM domains, such as mathematics and computer science, where the Routoo particularly excels.
-This impressive performance is largely attributed to the incorporation of specialized small models (around 7b parameters) that are fine-tuned for tasks in mathematics Yu et al.
-(2024); Shao et al.
-(2024) and coding Rozière et al.
-(2024); Guo et al.
-(2024) by the community.
+To further investigate the source of improvement in our models, we illustrates the distributions of performances on 17 sub-categories (Hendrycks et al., 2021) in Figure 3.
+A standout area of success is in STEM domains, such as mathematics and computer science, where Routoo particularly excels.
+This impressive performance is largely attributed to the incorporation of specialized small models (around 7b parameters) that are fine-tuned for tasks in mathematics (Yu et al., 2024; Shao et al., 2024) and coding (Rozière et al., 2024; Guo et al., 2024) by the community.
 Furthermore, this approach facilitates the identification of domains where there is a scarcity of experts.
@@
 Routing Distribution.
-Figure 6 presents the aggregated percentage usage of expert models by size for each Routoo pricing tier.
+Figure 4 presents the aggregated percentage usage of expert models by size for each Routoo pricing tier (\$ per 1M tokens).
 It reveals that higher-priced options tend to utilize larger models more frequently.
@@
 Finally, the cost of training a router is significantly lower than building foundational models e.g.
-GPT4 OpenAI et al.
-(2024) and Mixtral Jiang et al.
-(2024).
+GPT4 (OpenAI et al., 2024) and Mixtral (Jiang et al., 2024).
 This could pave a new path for building new frontiers at a much lower cost by integrating knowledge of current LLMs.
@@
 cost.
-We evaluated our architecture on MMLU benchmark Hendrycks et al.
-(2021), which is a MCQA dataset with 57 different domains coverage.
-Routoo (open-source) achieved competitive performance with Mixtral 8x7b model Jiang et al.
-(2024) with two-thirds of the inference cost.
+We evaluated our architecture on MMLU benchmark (Hendrycks et al., 2021), which is a MCQA dataset with 57 different domains coverage.
+Routoo (open-source) achieved competitive performance with Mixtral 8x7b model (Jiang et al., 2024) with two-thirds of the inference cost.
 Increasing the budget limitation allows Routoo (open-source) to outperform Mixtral model by 5% with the same level of cost.
-By integrating GPT4 OpenAI et al.
-(2024) model to underlying experts, our Routoo (mix) achieved competitive accuracy with GPT4 while reducing the inference cost by 50%, and even surpassing it while saving 25% of the cost.
+By integrating GPT4 (OpenAI et al., 2024) model to underlying LLMs, our Routoo (mix) achieved competitive accuracy with GPT4 while reducing the inference cost by 50%, and even surpassing it while saving 25% of the cost.
 In general, our Routoo models provide an efficient trade-off between cost and performance during the inference.
@@
 This insight is invaluable, enabling us to strategically develop domain-specific models where they are needed most.
-Limitations
-Due to GPU limitations, we fine-tuned our Routoo models using the LoRA technique Hu et al.
-(2021).
-For better performance, full fine-tuning methods could be employed to train our Routoo models.
-Our cost-aware decoder used a greedy technique to account for inference costs, though more comprehensive methods could better model the relationship between performance and cost.
-Resource constraints prevented us from using LLMs with 70 billion parameters as our underlying models.
-Incorporating these larger models, especially in the Routoo (mix) model, yields better performance compared to GPT4.
-Additionally, integrating other closed-source LLMs e.g.
-Claude  [^11] https://www.anthropic.com/news/claude-3-family and Gemini  [^12] https://deepmind.google/technologies/gemini/ variants into our Routoo models would increase the diversity of base LLMs.
+Furthermore, our architecture facilitates easy integration of additional optimization criteria, including cost, speed, and privacy considerations.
 Appendix A Optimal Routing of Open-source LLMs on MMLU
-In this experiment, we utilize all publicly available LLMs from the OpenLLM benchmark Beeching et al.
-(2023) that have fewer than 34 billion parameters.
+In this experiment, we utilize all publicly available LLMs from the OpenLLM benchmark (Beeching et al., 2023) that have fewer than 34 billion parameters.
 Given that each query is directed to the most effective and cost-efficient model by an ideal performance predictor module within Routoo, the upper bound for the Routoo model in this scenario is nearly 97.5%.
@@
 Model size Usage (%) 7b 66.4 13b 16.1 34b 17.5
-Table 2: The Distribution of usage among underlying LLMs with different sizes on MMLU Hendrycks et al. (2021) test set.
+Table 2: The Distribution of usage among underlying LLMs with different sizes on MMLU (Hendrycks et al., 2021) test set.
 The average size of the distribution above could be considered as an abstract model with approximately 13 billion parameters.
 Appendix B Synthetic Data Generation by GPT4
-The following input prompt is used for generating synthetic data with GPT4 model OpenAI et al. (2024):
-[…]
-[…]
-[…]
-[…]
-[…]
+The following input prompt is used for generating synthetic data with GPT4 model (OpenAI et al., 2024):
 […]
 For generation, we used OpenAI API  [^14] https://platform.openai.com/docs/overview.
-with $max_{tokens}=1000$ and $temperature=1$.
+with $\text{temperature}=0.7$.
+At each iteration, $N=100$, and we run it for 200 times.
 SUBJECT is used when the seed samples contain subject field e.g.
-development set of MMLU benchmark Hendrycks et al.
-(2021).
+development set of MMLU benchmark (Hendrycks et al., 2021).
 Seed samples are chosen randomly from datasets introduced in Section 4.2, alongside with development set of MMLU benchmark.